Write an equation in slope intercepth form of the line satisfying the given conditions. Please explain how you got your anwer. Thank You.
Parallel to the graph of 9x+3y=6 and passes through (5,3)

Answer by ilana(307)

OK, we want a line parallel to 9x+3y=6, so we need the slope of this. The easiest way to find this (for me) is to change the form of the equation to slope-intercept form, y=mx+b (m=slope). So, we get 3y=-9x+6, and then y=-3x+2, so the slope is -3. Parallel lines have the same slope, so we want a line with slope -3 through the point (5,3). We can use the point-slope form of a line, y-y1=m(x-x1). So we get our line, y-3=-3(x-5). We want slope intercept form again, so we can rewrite our line as y-3=-3x+15, so y=-3x+18, and we're done:).